he lengths of the sides of a triangle is extended ratio 4:7:8 the premeter is 76m. what are the lengths of the sides
Answer by ikleyn(52878)

.

The sides are 4x, 7x and 8x long, there x is their common measure.


Hence,  4x + 7x + 8x = 76 meters (the perimeter),  which implies


    19x = 76,

      x = 76/19 = 4 meters.


ANSWER.  The sides are 4x= 4*4 = 16 meters,  7x = 7*4 = 28 meters and  8x= 8*4 = 32 meters.

Solved.